# HOPE: A Reinforcement Learning-based Hybrid Policy Path Planner for Diverse Parking Scenarios

### v1

@@ Abstract
Path planning plays a pivotal role in automated parking, yet current methods struggle to efficiently handle the intricate and diverse parking scenarios. One potential solution is the reinforcement learning-based method, leveraging its exploration in unrecorded situations. However, a key challenge lies in training reinforcement learning methods is the inherent randomness in converging to a feasible policy. This paper introduces a novel solution, the Hybrid POlicy Path plannEr (HOPE), which integrates a reinforcement learning agent with Reeds-Shepp curves, enabling effective planning across diverse scenarios. The paper presents a method to calculate and implement an action mask mechanism in path planning, significantly boosting the efficiency and effectiveness of reinforcement learning training. A transformer is employed as the network structure to fuse environmental information and generate planned paths. To facilitate the training and evaluation of the proposed planner, we propose a criterion for categorizing the difficulty level of parking scenarios based on space and obstacle distribution. Experimental results demonstrate that our approach outperforms typical rule-based algorithms and traditional reinforcement learning methods, showcasing higher planning success rates and generalization across various scenarios. The code for our solution will be openly available on https://github.com/jiamiya/HOPE.
@@ I Introduction
Automatic parking technology has gained attention in recent years due to its potential benefits in safety and time efficiency [1].The entire automated parking systems comprise several vital components, including perception, planning, and control, among which path-planning algorithms play a crucial role [2]. The path-planning task in parking scenarios involves generating a feasible path that links the vehicle’s start position and the target position in the parking spots, as well as considering the vehicle’s physics constraints. While existing rule-based path-planning approaches prove practical and robust in most simple scenarios, their inherent difficulty in truly understanding the surrounding environment may lead to planning failures, especially when scenarios become more intricate [3].
@@
Learning-based planners have the potential to understand environments and intelligently plan routes through a data-driven approach, diverging from reliance on pre-defined planning methods rooted in human priors. While in various tasks, expert data is used as the ground truth for imitation learning, the scarcity of large-scale datasets for parking scenarios necessitates researchers to collect data manually [4]. This supervised learning approach poses the risk of overfitting the model to specific parking strategies due to the insufficient diversity in the training scenarios. Meanwhile, reinforcement learning has gathered increasing attention in the field of autonomous driving [5]. Indeed, through interaction with the environment, reinforcement learning enables the training of intelligent planners without the need for labeled trajectory ground truth. Nonetheless, training reinforcement learning agents in complex and diverse scenarios remains a challenging task [6]. It is straightforward for the agent to fit into fixed parking strategies but poses greater difficulty in obtaining generalization capability across various parking scenarios. Moreover, the agent faces challenges in conducting effective exploration, especially in complex scenarios with narrow parking spaces, which significantly impacts training efficiency.
@@
This paper focuses on employing reinforcement learning methods in the parking path planning task with static obstacles. To achieve efficient and effective learning under diverse parking scenarios, we propose reinforcement learning-based Hybrid pOlicy Path plannEr (HOPE). The hybrid policy planner is designed to leverage the reinforcement learning-based actor-critic method and a classical geometric-based method, the Reeds-Shepp (RS) curve [7]. A transformer-based structure is used as the information fusion network in actor and critic networks [8]. Meanwhile, since the diversity of scenario difficulty has a significant impact on training and testing, we rank the difficulty of scenarios by referencing related standards for automated parking. Overall, the main contributions of this paper include:
  • We develop a hybrid policy method for parking path planning, which takes advantage of both learning-based and rule-based methods to achieve generalization capability under diverse parking scenarios.
@@ II-A Non-learning-based parking path-planning
Non-learning-based path planning methods under parking scenarios mainly include geometric-based methods and sampling-and-search-based methods [9]. Geometric planners construct paths connecting the start position and the target position using different kinds of geometric curves [10, 11]. The most classic geometric methods include the Dubins curves and the RS curves [12, 7]. Building upon the RS curve, several improvements have been made to enhance its performance in parking scenarios [13, 14]. The sampling-and-search-based approach discretizes and searches either in the state space or in the control space to find a valid path. One widely employed approach is the Hybrid A* method [15], which devises a variant of the A* algorithm that incorporates the kinematic state space of the vehicle to derive a kinematically feasible trajectory. Hybrid A* was initially utilized in the Defense Advanced Research Projects Agency Challenge (DARPA) and has since undergone improvements and application to path planning tasks in parking scenarios [16, 17]. Both geometric and sampling-and-search-based methods, as rule-based methods, leverage human prior knowledge to design algorithms. In most common cases, these priors can serve as a strong fallback to obtain satisfactory solutions. However, it remains challenging for rule-based methods to achieve human-level proficiency in complex and diverse parking scenarios [18].
Another category of methods involving path planning is the optimization-based trajectory planning approach. These methods solve the trajectory planning problem by formulating it as an optimal control problem. Related works have made improvements in collision avoidance constraint formulation [19], iterative solution efficiency [9], and robustness [20]. While this approach can yield smoother paths that adhere to vehicle kinematics compared to rule-based path-planning algorithms, the optimization process relies on another path-planning algorithm, often the Hybrid A* method [21, 19, 20], to obtain initial solutions. Therefore, these methods are typically used in conjunction with path planning algorithms to generate smoother trajectories rather than being employed independently for feasible path generation.
@@ II-B Learning-based parking path-planning
Learning-based methods represent a potential avenue to enhance the planner’s performance in the task of parking. Existing methods mainly include imitating learning-based and reinforcement learning-based methods. Imitating learning requires the learnable planner to fit on the ground truth data. Liu et al. employed a neural network to align parking paths more closely with human behavior [22]. Rathour S et al. derived a parking policy by imitating trajectories collected from expert drivers [23]. Other works include implementing the deep neural network or deep recurrent neural network on real vehicles [24, 25]. However, existing imitation learning-based approaches still lack diversity and complexity in scenarios, and the behavior cloners cannot surpass the performance of their imitation targets [26].
In the reinforcement learning-based approaches, intelligent planners can be obtained without the labeled ground truth by interacting with the environment. The Monte Carlo tree search method was applied to search the available path in parallel parking scenarios [27]. Bernhard J et al. learned the heuristics with Deep Q-learning (DQL) and improved the path search process in the Hybrid A* algorithm [28, 29]. Du et al. directly employed DQL to train an agent to produce single-step path planning results iteratively in a parallel and vertical case [30]. Yuan et al. proposed a hierarchical planning approach where high-level reinforcement learning agents are trained to generate initial reference solutions [31]. The existing works typically map the elements of the task to the components of reinforcement learning and directly employ them to train the agent in a limited number of scenarios. As more complex and diverse parking scenarios are introduced, it would be harder for reinforcement learning-based methods to explore the proper parking policy, making it challenging to obtain planners with generalization capabilities.
@@ III-A Reinforcement learning for path planning
The reinforcement learning problem can be addressed as policy learning in a Markov decision process(MDP) defined by a 4-element tuple $(\mathcal{S},\mathcal{A},p,r)$, where $\mathcal{S}$ is the state space, and $\mathcal{A}$ is the action space. In the path-planning problem, $s_{t}\in\mathcal{S}$ includes the vehicle position and orientation $p_{t}=(x_{t},y_{t},\theta_{t})$ as well as other observable information about obstacles and target parking space. We selecte velocity $v$ and steering angle $\delta$ as the action space $\mathcal{A}=\{a=(v,\delta)\}$. $p$ denotes the state transition probability density of the next state $s_{t+1}$ given the current state $s_{t}$ and action $a_{t}$. In practice, the transition is modeled using the single-track bicycle model, as implemented in [32], and this uncertainty is not considered in a deterministic environment. A reward $r=r(s_{t+1},a_{t})$ is given by the environment $E$ based on the state and action in each step. The objective is to learn the policy $\pi(a_{t}|s_{t})$ that maximizes the future reward $R_{t}=\sum_{i=t}^{T}\gamma^{(i-t)}r(s_{i},a_{i})$ with a discounting factor $\gamma\in[0,1]$. Given the start point $p_{0}$ and the target $p_{T}$, the feasible parking path $P=\{p_{0},p_{1},p_{2},...,p_{T}\}$ can be obtained iteratively using the policy $\pi$.
@@ III-B Reeds-Shepp curve
RS curve is designed to generate the shortest path that links two positions with the minimal steering radius of the vehicle and straight lines. It is mathematically proved that the shortest path belongs to 48 types of curves and can be represented by one of the following 9 expressions:
@@ IV-A Overview of architecture
The overall reinforcement learning-based path planning framework and network structure are shown in Figure 1. To enhance training efficiency, we employ a hybrid policy reinforcement learning approach. Our proposed HOPE combines the learnable policy from original reinforcement learning with a rule-based policy derived from the RS curve. In each step, the agent outputs the action, the single-step path planning result, based on the current state given by the environment. The action is then adjusted using the action mask mechanism before interacting with the environment. We employ four forms of network input as state representation, including information about the distance to obstacles, target parking space, drivable action space, and historical trajectory, encoded in either vectors or images. A transformer-based structure with learnable view encoding is designed to fuse the inputs and get the outputs in the actor and critic network [35]. We also utilize the auto-encoder structure to pre-train the image encoder. More details of our network structure and reward function are discussed in the appendix. The pseudo-code of our method is shown in Algorithm 1.
@@ IV-B Hybrid policy
To improve the exploring and training efficiency, we combine the RL policy $\pi_{\theta}$ and the derived RS policy $\pi_{RS}$ to facilitate the planning process. The hybrid process at timestamp $t$ can be expressed as a function $h_{t}:\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ that maps the actions $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ from RL policy and $a^{\prime}_{t}\sim\pi_{RS}(\cdot|s_{t})$ from RS policy to the hybridized action $\tilde{a}_{t}=h_{t}(a_{t},a^{\prime}_{t})\sim\pi_{h}(\cdot|s_{t})$. Such an action hybrid process can be regarded as the procedure of making a suitable choice between two actions. Specifically, it activates the rule-based policy at certain timestamps in addition to the reinforcement learning policy, thereby enhancing the robustness of the overall algorithm.
@@ IV-B1 Rule-based policy from Reeds-Shepp curve
  • We add the computation of the paths in expression $S(|)C(|)S$ as these paths, though strictly suboptimal in the absence of obstacles, may be feasible when obstacles are considered. Both real-world parking experience and experiments suggest this is effective for vertical parking.
@@ IV-B2 Exploring and learning with RS policy
During training, the agent explores and interacts with the environment using the hybridized action $\tilde{a}_{t}$ rather than raw action $a_{t}$. This choice also affects the updates on the policy $\pi_{\theta}$ and Q function or value function. For the Bellman equation in Equation 2, since the states and actions are obtained with the hybrid policy $\pi_{h}$, the update on the parameter $V_{\psi}$ and $Q_{\phi}$ can be written as:
@@
Therefore, the update process shown in Equation 6 has the same convergency as Equation 2. For policy updates in PPO, the probability ratio $r_{t}(\theta)$ in Equation 3 can be rewritten as:
@@
Generally, gradients on the Q function and value function are modified by using the data collected by $\pi_{h}$ instead of $\pi_{\theta}$, since they are now estimations on distribution $\rho_{\pi_{h}}$. We still update the original RL policy $\pi_{\theta}$ as shown in Equation 7 and 8, which indicates that the hybrid policy $\pi_{h}$, as well as the RL policy $\pi_{\theta}$, are optimized. Thus, better performance can be only achieved when $Q_{\pi_{h}}\geq Q_{\pi_{\theta}}$ and $V_{\pi_{h}}\geq V_{\pi_{\theta}}$. In practice, the RS policy is activated only when: 1) the distance from the vehicle’s position to the target position is smaller than a threshold $d_{rs}$, and 2) there exists a collision-free RS curve from the vehicle’s position to the target position. By the hybrid policy approach, during the initial phases of the learning process when the agent is not well-trained, the RS method serves as an alternative strategy to offer additional positive examples for updates. The agent can thus learn how to adjust the vehicle’s pose to explore feasible parking routes.
@@ IV-C Action mask
In some discrete-space reinforcement learning tasks, employing action masks to filter out invalid actions can enhance training efficiency [36]. However, in path planning tasks, the computation of an action mask, which includes collision detection for all feasible actions of the agent vehicle across diverse state spaces, is computationally expensive. Here, we present a method for calculating and utilizing the action mask to enhance the training efficiency of reinforcement learning in path-planning tasks. Specifically, we use $Collide(s_{t})$ to denote the event whether the vehicle at state $s_{t}$ collides with obstacles. The action mask provides information about the largest safe step velocity $v^{*}$ at any given steering angle $\delta$:

### v2

@@ Abstract
Automated parking stands as a highly anticipated application of autonomous driving technology. However, existing path planning methodologies fall short of addressing this need due to their incapability to handle the diverse and complex parking scenarios in reality. While non-learning methods provide reliable planning results, they are vulnerable to intricate occasions, whereas learning-based ones are good at exploration but unstable in converging to feasible solutions. To leverage the strengths of both approaches, we introduce Hybrid pOlicy Path plannEr (HOPE). This novel solution integrates a reinforcement learning agent with Reeds-Shepp curves, enabling effective planning across diverse scenarios. HOPE guides the exploration of the reinforcement learning agent by applying an action mask mechanism and employs a transformer to integrate the perceived environmental information with the mask. To facilitate the training and evaluation of the proposed planner, we propose a criterion for categorizing the difficulty level of parking scenarios based on space and obstacle distribution. Experimental results demonstrate that our approach outperforms typical rule-based algorithms and traditional reinforcement learning methods, showing higher planning success rates and generalization across various scenarios. We also conduct real-world experiments to verify the practicability of HOPE. The code for our solution will be openly available on https://github.com/jiamiya/HOPE.
@@ I Introduction
Automated parking is a tempting technology to improve driving safety and efficiency [1]. An automated parking system comprises several vital components, including perception, planning, and control, in which path-planning algorithms play a crucial role [2]. In parking scenarios, the path-planning task involves generating a feasible path from the start position to the target parking spots under specific physics constraints. Compared to other scenarios, finding a feasible path in parking scenarios is usually more challenging because of the lower error tolerance of the target spot and the lack of navigational reference lines [3]. Additionally, the limited space and surrounding obstacles shrink the number of potential solutions. While existing path-planning approaches have proved practical and robust in most simple scenarios, their inherent difficulty in truly understanding the surrounding environment may lead to planning failures, especially when scenarios become more intricate [4].
@@
Learning-based planners have the potential to understand environments and intelligently plan routes through a data-driven approach, diverging from reliance on pre-defined planning methods rooted in human priors. While expert data is used as the ground truth for imitation learning in various tasks, the scarcity of large-scale datasets for parking scenarios necessitates researchers to collect data manually [5]. This supervised learning approach risks overfitting the model to specific parking strategies due to the insufficient diversity in the training scenarios. Meanwhile, Reinforcement Learning (RL) has gathered increasing attention in the field of autonomous driving [6]. Indeed, through interaction with the environment, RL methods enable the training of agents without labeling trajectory ground truth. Nonetheless, training RL agents in complex and diverse scenarios remains a challenging task [7]. It is straightforward for the agent to fit into fixed parking strategies but poses greater difficulty in obtaining generalization capability across various parking scenarios. Moreover, the agent faces challenges in conducting effective exploration, especially in complex scenarios with narrow parking spaces, significantly impacting the training efficiency.
@@
This paper focuses on employing RL methods in the parking path planning task with static obstacles. To achieve efficient and effective learning under diverse parking scenarios, we propose reinforcement learning-based Hybrid pOlicy Path plannEr (HOPE). The hybrid policy planner is designed to leverage RL-based methods and a classical geometric-based path planning method, the Reeds-Shepp (RS) curve [8]. A transformer-based structure is used as the information fusion network in actor and critic networks [9]. Since the diversity of scenario difficulty has a significant impact on training and testing, we rank the difficulty of scenarios by referencing related standards for automated parking. Overall, the main contributions of this paper include:
  • We develop a hybrid policy method for parking path planning which achieves over 97% success rate across diverse and challenging parking scenarios and verify its generalization ability with real-world experiments.
@@ II-A Non-learning-based parking path-planning
Non-learning-based path planning methods under parking scenarios mainly include geometric-based methods and sampling-and-search-based methods [3]. Geometric planners construct paths connecting the start position and the target position using different kinds of geometric curves [10, 11]. Typical geometric methods include the Dubins curves and the RS curves [12, 8]. Building upon the RS curve, several improvements have been made to enhance its performance in parking scenarios [13, 14]. The sampling-and-search-based approach discretizes and searches either in the state space or in the control space to find a valid path. One widely employed approach is the Hybrid A* method [15], which devises a variant of the A* algorithm that incorporates the kinematic state space of the vehicle to derive a kinematically feasible trajectory. Hybrid A* was initially utilized in the Defense Advanced Research Projects Agency Challenge (DARPA) and has since undergone improvements and application to path planning tasks in parking scenarios [16, 17]. Both geometric and sampling-and-search-based methods, as rule-based methods, leverage human prior knowledge to design algorithms. In most common cases, these priors can serve as a strong fallback to obtain satisfactory solutions. However, these methods can hardly achieve human-level proficiency in complex and diverse parking scenarios [18].
Another category of methods involving path planning is the optimization-based trajectory planning approach. These methods solve the trajectory planning problem by formulating it as an optimal control problem. Related works have made improvements in collision avoidance constraint formulation [19], iterative solution efficiency [3], and robustness [20]. While this approach can produce smoother paths adhering to vehicle kinematics, the optimization process relies on another path-planning algorithm, often the Hybrid A* method [21, 19, 20], to obtain initial solutions. This reliance helps improve computational efficiency and allows these methods to focus on optimizing smoother trajectories, as well as determining the limit of planning success rate.
@@ II-B Learning-based parking path-planning
Learning-based methods represent a potential avenue to enhance the planner’s performance in the task of parking. Existing methods mainly include imitating learning-based and reinforcement learning-based methods. Imitating learning requires the learnable planner to fit on the ground truth data. Liu et al. employed a neural network to align parking paths more closely with human behavior [22]. Rathour S et al. derived a parking policy by imitating trajectories collected from expert drivers [23]. Other works include implementing the deep neural network or deep recurrent neural network as the behavior cloner on real vehicles [24, 25]. However, imitation learning-based methods are not good at handling the scenarios beyond their training sets, and the behavior cloners cannot outperform their imitation targets in terms of performance [26].
In RL-based approaches, the agents are developed without any labeled ground truth. The Monte Carlo tree search method was applied to search the available path in parallel parking scenarios [27]. Bernhard J et al. learned the heuristics with Deep Q-learning (DQL) and improved the path search process in the Hybrid A* algorithm [28, 29]. Du et al. directly employed DQL to train an agent to produce single-step path planning results iteratively in a parallel and vertical case [30]. Yuan et al. proposed a hierarchical planning approach where high-level RL agents are trained to generate initial reference solutions [31]. The existing works typically map the elements of the task to the components of RL and directly employ them to train the agent in a limited number of scenarios. As more complex and diverse parking scenarios are introduced, it would be harder for RL-based methods to explore the proper parking policy, making it challenging to obtain planners with generalization capabilities.
@@ III-A Reinforcement learning for path planning
The RL problem can be addressed as policy learning in a Markov decision process (MDP) defined by a 4-element tuple $(\mathcal{S},\mathcal{A},p,r)$, where $\mathcal{S}$ is the state space, and $\mathcal{A}$ is the action space. In the path-planning problem, $s_{t}\in\mathcal{S}$ includes the vehicle position and orientation $p_{t}=(x_{t},y_{t},\theta_{t})$ as well as other observable information about obstacles and target parking space. We selecte velocity $v$ and steering angle $\delta$ as the action space $\mathcal{A}=\{a=(v,\delta)\}$. $p$ denotes the state transition probability density of the next state $s_{t+1}$ given the current state $s_{t}$ and action $a_{t}$. In practice, the transition is modeled using the single-track bicycle model, as implemented in [32], and this uncertainty is not considered in a deterministic environment. A reward $r=r(s_{t+1},a_{t})$ is given by the environment $E$ based on the state and action in each interaction step. The objective is to learn the policy $\pi(a_{t}|s_{t})$ that maximizes the future reward $R_{t}=\sum_{i=t}^{T}\gamma^{(i-t)}r(s_{i},a_{i})$ with a discounting factor $\gamma\in[0,1]$. Given the start point $p_{0}$ and the target $p_{T}$, the feasible parking path $P=\{p_{0},p_{1},p_{2},...,p_{T}\}$ can be obtained iteratively using the policy $\pi$.
@@ III-B Reeds-Shepp curve
RS curve is designed to generate the shortest path that links two positions with the minimal steering radius of the vehicle and straight lines. It is mathematically proved that the shortest path belongs to 48 types of curves and can be represented by one of the following 9 expressions  [^1] Visualization examples of RS curves can be found in https://tactics2d.readthedocs.io/en/latest/tutorial/interpolator_visualization/:
@@ IV-A Overview of architecture
The overall RL-based path planning framework and network structure are shown in Figure 1. To enhance training efficiency, we employ a hybrid policy reinforcement learning approach. Our proposed HOPE combines the learnable policy from original reinforcement learning with a rule-based policy derived from the RS curve. In each interaction step, the agent outputs the action, namely the single-step path planning result, based on the current state given by the environment. The action is then adjusted using the action mask mechanism before interacting with the environment. We employ four forms of network input as state representation, including information about the distance to obstacles, target parking space, drivable action space, and historical trajectory, encoded in either vectors or images. A transformer-based structure with learnable view encoding is designed to fuse the inputs and get the outputs in the actor and critic network [35]. We also utilize the auto-encoder structure to pre-train the image encoder. More details of our input modality and reward function are discussed in the appendix A and C. The pseudo-code of our method is shown in Algorithm 1.
@@ IV-B Hybrid policy
To improve the exploring and training efficiency, we combine the RL policy $\pi_{\theta}$ and the derived RS policy $\pi_{RS}$ to facilitate the planning process. The hybrid process at timestamp $t$ can be expressed as a function $h_{t}:\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ that maps the actions $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ from RL policy and $a^{\prime}_{t}\sim\pi_{RS}(\cdot|s_{t})$ from RS policy to the hybridized action $\tilde{a}_{t}=h_{t}(a_{t},a^{\prime}_{t})\sim\pi_{h}(\cdot|s_{t})$. Such an action hybrid process can be regarded as the procedure of making a suitable switch between two actions. Specifically, it activates the rule-based policy at certain timestamps and otherwise the reinforcement learning policy is used.
@@ IV-B1 Rule-based policy from Reeds-Shepp curve
  • We add the computation of the paths in expression $S(|)C(|)S$ as these paths, though strictly suboptimal in the absence of obstacles, may be feasible when obstacles are considered. Experiments in appendix B suggest this is effective for vertical parking.
@@ IV-B2 Exploring and learning with RS policy
During training, the agent explores and interacts with the environment using the hybridized action $\tilde{a}_{t}$ rather than raw action $a_{t}$. This choice also affects the updates on the policy $\pi_{\theta}$ and Q function or value function. For the Bellman equation in Equation 2, the update on the parameter $V_{\psi}$ and $Q_{\phi}$ can be written as:
@@
Since the convergence of the Bellman equation for updating $V_{\psi}$ and $Q_{\phi}$ is independent of the specific policy, when we replace the policy $\pi$ with $\pi_{h}$, the iterative updates for $V_{\psi}$ and $Q_{\phi}$ still converge. For policy updates in PPO, the probability ratio $r_{t}(\theta)$ in Equation 3 can be rewritten as:
@@
Generally, gradients on the Q function and value function are modified by using the data collected by $\pi_{h}$ instead of $\pi_{\theta}$, since they are now estimations on distribution $\rho_{\pi_{h}}$. We still update the original RL policy $\pi_{\theta}$ as shown in Equation 7 and 8, which indicates that the hybrid policy $\pi_{h}$, as well as the RL policy $\pi_{\theta}$, are optimized. In practice, we apply a switch strategy that the RS policy is activated only when: 1) the distance from the vehicle’s position to the target position is smaller than a threshold $d_{rs}$, and 2) there exists a collision-free RS curve from the vehicle’s position to the target position. By the hybrid policy approach, during the initial phases of the learning process when the agent is not well-trained, the RS method serves as an alternative strategy to offer additional positive examples for updates. The agent can thus learn how to adjust the vehicle’s pose to explore feasible parking routes.
@@ IV-C Action mask
Action masks have been employed in some discrete-space reinforcement learning tasks to filter out invalid actions, which can enhance training efficiency and ensure alignment with deployment conditions [36]. By incorporating action masks, agents can focus on making complex decisions without spending extensive time learning calculable constraints. Additionally, this approach helps prevent invalid behaviors in applications. However, in path planning tasks, the computation of an action mask, which includes collision detection for all feasible actions of the agent vehicle across diverse state spaces, is computationally expensive. Here, we present a method for calculating and utilizing the action mask to enhance the training efficiency of reinforcement learning in path-planning tasks. Specifically, we use $Collide(s_{t})$ to denote the event whether the vehicle at state $s_{t}$ collides with obstacles. The action mask provides information about the largest safe step velocity $v^{*}$ at any given steering angle $\delta$: